OpenStudy (anonymous):

5.0x 10 24 molecules equals how many liters

OpenStudy (chmvijay):

6.023*10^23 molecules =22.4 litres 5.0*10^24 molecules = 22.4*5.0*10^24 /6.023*10^23

OpenStudy (chmvijay):

WELCOME TO OPEN STUDY